a) The weekly wages of employees of Volta gold are normally distributed about a mean of $1,250 with a standard deviation of $120. Find the probability of an employee having a weekly wage lying;
i) Between $1,320 and $970
ii) Under $1,400
iii) Over $1,290

b) An achievement test was administered to a class of 20,000 students. The mean score was 80 and the standard deviation was 11. If Lingard scored 72 in the test, how many students did better than him?
1
Expert's answer
2021-07-19T14:13:02-0400

a)

"\mu=1250 \\

\sigma=120 \\

i. \; P(970<X<1320) = P(X<1320) -P(X<970) \\

= P(Z< \frac{1320-1250}{120}) -P(Z< \frac{970-1250}{120}) \\

= P(Z< 0.5833) -P(Z< -2.333) \\

= 0.72005-0.00982 \\

= 0.71023 \\

ii. \;P(X<1400) = P(Z< \frac{1400-1250}{120}) \\

=P(Z< 1.25) \\

= 0.8943 \\

iii. \; P(X>1290) = 1 -P(X<1290) \\

= 1 -P(Z< \frac{1290-1250}{120}) \\

= 1 -P(Z<0.3333) \\

= 1 -0.63043 \\

= 0.36957"

b)

"N=20000 \\

\mu=80 \\

\sigma=11 \\

P(X>72) = 1 -P(X<72) \\

= 1 -P(Z< \frac{72-80}{11}) \\

= 1 -P(Z < -0.7272) \\

= 1 -0.2336 \\

= 0.7664 \\

n(X>72) = 20000 \times 0.7664 = 15328"

Answer: 15328 students were better than Lingard